# Post-Training Quantization of OpenPangu Models for Efficient Deployment on Atlas A2

### v1

@@ Abstract
Huawei’s openPangu-Embedded-1B and openPangu-Embedded-7B, variants of the openPangu large language model, integrate three distinct Chain-of-Thought (CoT) reasoning paradigms, namely slow_think, auto_think, and no_think. While these CoT modes enhance reasoning capabilities, their generation of extended reasoning traces introduces substantial memory and latency overheads, posing challenges for practical deployment on Ascend NPUs. This paper addresses these computational constraints by leveraging low-bit quantization, which transforms FP16 computations into more efficient integer arithmetic. We introduce a unified low-bit inference framework, supporting INT8 (W8A8) and W4A8 quantization, specifically optimized for openPangu-Embedded models on the Atlas A2. Our comprehensive evaluation, conducted across all three CoT modes on code generation benchmarks (HumanEval and MBPP), demonstrates the efficacy of this approach. INT8 quantization consistently preserves over 90% of the FP16 baseline accuracy and achieves a 1.5x prefill speedup on the Atlas A2. Furthermore, W4A8 quantization significantly reduces memory consumption, albeit with a moderate trade-off in accuracy. These findings collectively indicate that low-bit quantization effectively facilitates efficient CoT reasoning on Ascend NPUs, maintaining high model fidelity.
@@ 1 Introduction
Large language models are increasingly being deployed on edge devices to facilitate on-device reasoning and code generation (zheng2025reviewedgelargelanguage; Girija_2025). Recently, Huawei introduced the openPangu-Embedded series, which includes the 7B and 1B variants (chen2025panguembeddedefficientdualsystem). These models support three Chain-of-Thought (CoT) reasoning paradigms including slow_think, auto_think, and no_think (wei2023chainofthoughtpromptingelicitsreasoning; zhang2022automaticchainthoughtprompting). Each of these modes varies in how intermediate reasoning steps are generated, influencing both latency and output quality. Despite the strong performance of these models, their original FP16 versions require substantial memory and computational resources, which limit their practical deployment on NPU accelerators (xiao2024smoothquantaccurateefficientposttraining; ashkboos2024quarot; zhao2024atomlowbitquantizationefficient), such as the Huawei Atlas A2, where memory bandwidth and storage capacity present significant constraints.
@@
To address the high memory and computational demands of FP16 inference, we apply quantization to the openPangu-Embedded models. This technique replaces floating-point parameters and activations with low-bit integer representations, significantly reducing both model size and computational complexity while aiming to maintain task performance (wei2023outliersuppressionaccuratequantization; yao2023zeroquantv2exploringposttrainingquantization; liu2024crossquantposttrainingquantizationmethod). Specifically, we focus on two quantization configurations commonly used for efficient large language model deployment, namely INT8 (W8A8) and W4A8 (dettmers2022llmint88bitmatrixmultiplication; lee2024enhancingcomputationefficiencylarge). Both configurations leverage the integer arithmetic capabilities of Ascend NPUs, offering a practical trade-off between efficiency and accuracy for on-device applications.
@@
However, most existing quantization techniques are designed and optimized for GPU architectures and are not easily transferable to Ascend NPUs(wang2019haqhardwareawareautomatedquantization). Adapting quantized models to the Atlas A2 presents significant challenges due to its unique compute primitives, memory hierarchy, and integer execution model. Efficient low-bit inference on this architecture requires tight integration between the quantization format, memory layout, and compute primitives to minimize redundant data movement and avoid underutilizing hardware resources (feng2022tensorirabstractionautomatictensorized; frantar2023gptqaccurateposttrainingquantization). To facilitate practical evaluation, we develop hardware-aware quantized inference paths that are fully optimized for the Atlas A2 execution environment.
@@
Beyond system-level deployment, we observe that quantization impacts the reasoning behavior of the openPangu-Embedded models. The three CoT paradigms slow_think, auto_think, and no_think exhibit distinct characteristics and respond differently to low-bit conversion(chen2025panguembeddedefficientdualsystem). In slow_think, the model generates detailed intermediate steps to support complex reasoning. In no_think, the reasoning process is more condensed, retaining the stepwise structure but operating with higher efficiency. The auto_think mode adaptively switches between these two strategies based on input characteristics. Our evaluation reveals that quantization affects the behavior of all three modes, prompting a systematic investigation into the impact of quantization on both model execution and reasoning quality.
@@
(3) Analysis of how quantization affects CoT reasoning. We analyze the impact of INT8 quantization on CoT reasoning behavior in openPangu-Embedded-1B and 7B across the three reasoning modes, using HumanEval and MBPP as evaluation benchmarks. Our findings show that the transition from FP16 to INT8 introduces minimal disruption to the reasoning process in most scenarios, indicating that quantized models can still effectively perform structured multi-step inference.
@@ 2 Preliminary
To enable efficient inference of large language models on resource-constrained hardware, quantization techniques are used to approximate high-precision floating-point values with low-bit integer representations. This approach reduces both memory footprint and computational cost, while aiming to preserve model accuracy (dettmers2022llmint88bitmatrixmultiplication; yao2023zeroquantv2exploringposttrainingquantization; jin2024comprehensiveevaluationquantizationstrategies).
@@
The fidelity of quantization is highly dependent on the granularity of the scaling factor(nagel2021whitepaperneuralnetwork). Per-tensor quantization uses a single scale for an entire tensor, offering memory efficiency but often resulting in lower accuracy(bondarenko2021understandingovercomingchallengesefficient). Per-channel quantization assigns a separate scale to each output channel of weight matrices and is widely used due to its effectiveness (lin2023awq). For activations, per-token quantization computes one scale per token across its entire feature dimension (shao2024omniquantomnidirectionallycalibratedquantization). To balance flexibility and overhead, group-wise quantization divides each tensor into fixed-size groups, each with its own scale (dettmers2023qloraefficientfinetuningquantized; zhao2024atomlowbitquantizationefficient). These strategies are essential for maintaining accuracy under low-bit constraints.
@@ 3.2 Quantization Configurations and Evaluation Protocol
The first variant adopts SmoothQuant (xiao2024smoothquantaccurateefficientposttraining), which balances the quantization difficulty between activations and weights using a diagonal scaling ${S}=\text{diag}({s})$. The linear transformation is rewritten as:
@@
The second variant employs Hadamard rotation (ashkboos2024quarot; liu2024spinquantllmquantizationlearned), which applies an orthogonal transform to the weight matrix before quantization. The output is computed as:
@@
where H{H} is a normalized Hadamard matrix. This rotation preserves the mathematical equivalence in full precision while promoting a more uniform weight distribution, which is better suited for low-bit quantization.
@@ 4.1 Experimental Setup
Quantization Configuration. We apply post-training quantization without retraining, using calibrated scales derived from downstream task data. All quantized models employ symmetric quantization. Our evaluation covers INT8 and three W4A8 configurations, including a baseline W4A8 setting as well as two enhanced variants that incorporate SmoothQuant(xiao2024smoothquantaccurateefficientposttraining) and Hadamard rotation(ashkboos2024quarot) respectively. For performance measurement on the Atlas A2, all configurations are deployed using hardware-optimized GEMM interfaces tailored to their corresponding bitwidths. Accuracy evaluation is conducted under a unified quantization pipeline to ensure consistent calibration and quantization procedures across all settings.
Benchmarks. We evaluate model performance on two standard code generation benchmarks, HumanEval (human-eval) and MBPP (mbpp). For each benchmark, we test all three CoT reasoning paradigms supported by openPangu-Embedded, including slow_think, auto_think, and no_think. These modes are enabled at inference time by appending the corresponding directive to the input prompt. We compare INT8 quantization against the FP16 baseline to assess accuracy retention and inference efficiency, and further analyze how the evaluated W4A8 configurations influence model fidelity under a fixed 4-bit weight setting on Ascend NPUs.
@@ 4.2 Main Results
Table 1 reports the accuracy on HumanEval and MBPP for openPangu-Embedded models under the no_think, auto_think, and slow_think modes, with FP16 and INT8 quantization. For the 7B model, INT8 matches or exceeds the performance of FP16 across all settings. In no_think, the scores are similar (HumanEval: 85.37 vs 85.37; MBPP: 77.04 vs 78.21). In auto_think, INT8 improves by 2.44 on HumanEval and 3.11 on MBPP. Under slow_think, INT8 achieves 95.73 on HumanEval and 79.77 on MBPP, compared to FP16’s 95.12 and 77.43, respectively.
For the 1B model, INT8 shows a modest drop in no_think (HumanEval: 66.46 vs 70.73), but gains in auto_think (HumanEval: 70.12 vs 67.07; MBPP: 65.37 vs 60.70). Under slow_think, INT8 again matches or surpasses FP16 (HumanEval: 66.46 vs 65.24; MBPP: 62.26 vs 61.87). These results confirm that INT8 quantization preserves high code generation accuracy across model scales and reasoning strategies, supporting efficient deployment without significant loss in fidelity.
@@
1B no_think FP16 70.73 58.75
INT8 66.46 58.37
auto_think FP16 67.07 60.70
INT8 70.12 65.37
slow_think FP16 65.24 61.87
@@ 4.4 CoT Analysis
We analyze the impact of quantization on CoT output length across the openPangu-Embedded-1B and 7B models on HumanEval and MBPP. As shown in Figure 2, quantization from FP16 to INT8 has little effect on output length in most configurations. Notably, the 7B model consistently produces shorter CoT traces than the 1B model across all reasoning modes and benchmarks, suggesting that larger models may adopt more concise reasoning strategies.
Figure 2: Average word count on HumanEval and MBPP for FP16 and INT8 openPangu-Embedded-1B/7B under three CoT modes. Quantization has only a limited effect on output length in most cases, while the 7B model consistently produces shorter traces than the 1B model, hinting at more compact reasoning at scale.
@@
Beyond surface-level variations in output length and phrasing, we further examine the structural integrity of CoT reasoning under quantization. Specifically, we identify and analyze a critical degradation pattern termed repetitive generation, where model outputs terminate with identical phrases repeated multiple times. Figure 4 quantifies this phenomenon across configurations on HumanEval. The 1B FP16 model exhibits high susceptibility to repetitive outputs, particularly in slow_think mode (34.15%) and auto_think mode (25.00%). INT8 quantization significantly suppresses this behavior in 1B models, reducing repetitive occurrences by 12.2 percentage points in slow_think. In contrast, the 7B model demonstrates inherent robustness, with repetitive ratios consistently below 2.5% regardless of precision. Crucially, repetitive generation strongly correlates with functional failure, as non-repetitive samples achieve an average accuracy of 87.39%, while repetitive samples achieve only 18.24%. This indicates that repetitive generation fundamentally disrupts reasoning integrity.
Figure 4: Frequency of repetitive generation patterns across reasoning modes and model configurations on HumanEval. Repetitive generation is defined as terminal output segments containing identical phrases repeated until sequence termination. Bars represent the percentage of test samples exhibiting this behavior under FP16 and INT8 precision for 1B and 7B model variants.
@@
Post-training quantization has become a prominent approach for compressing large language models without requiring access to original training data. It is broadly classified into weight-only quantization and weight–activation quantization. Weight-only approaches (lin2023awq; frantar2023gptqaccurateposttrainingquantization; kim2024squeezellmdenseandsparsequantization) compress model weights into low-bit representations while restoring them to high precision (e.g., FP16) during GEMM computations. Although this strategy alleviates memory bandwidth demands, the underlying arithmetic operations remain in high precision, which continues to impose a substantial bottleneck on inference efficiency (lin2025qservew4a8kv4quantizationcodesign). Consequently, significant opportunities remain for accelerating large language model inference. In contrast, weight–activation methods (yao2022zeroquantefficientaffordableposttraining; lee2024owqoutlierawareweightquantization) quantize both weights and activations into low-bit formats, enabling GEMM to be performed entirely in low precision. This paradigm mitigates both memory bandwidth and computational constraints but frequently incurs pronounced accuracy degradation due to outlier activations. To address this issue, mathematically equivalent transformation techniques (xiao2024smoothquantaccurateefficientposttraining; shao2024omniquantomnidirectionallycalibratedquantization) employ channel-wise smoothing strategies. By transferring activation outliers into the weight domain, these methods effectively suppress quantization errors. More recently, rotation-based weight–activation methods (lin2024duquant; liu2024spinquantllmquantizationlearned; ashkboos2024quarot) have been proposed and demonstrated remarkable success in maintaining model accuracy even at 4-bit precision. Complementing uniform low-bit schemes, mixed-precision quantization achieves a better accuracy-efficiency trade-off by assigning higher bitwidths to outlier-prone components and lower bitwidths to stable ones (zhao2024atomlowbitquantizationefficient; liu2025micromixefficientmixedprecisionquantization).
Impact of quantization on reasoning behavior. Recent studies have shown that quantization can affect the reasoning behavior of large language models beyond mere task accuracy. Empirical evidence suggests that quantization error may accumulate over CoT steps, disproportionately degrading performance on complex reasoning tasks (liu2025quantizationhurtsreasoningempirical). Furthermore, while moderate quantization typically preserves output length, which serves as a proxy for reasoning depth, aggressive low-bit schemes can lead to unstable generation and disrupted logical flow (li2025quantizationmeetsreasoningexploring). However, these analyses are primarily focused on general benchmarks and NVIDIA platforms, with limited attention given to structured reasoning modes like slow_think or auto_think on specialized hardware, such as Ascend NPUs.

### v2

@@ Abstract
Huawei’s openPangu-Embedded-1B and openPangu-Embedded-7B are variants of the openPangu large language model, designed for efficient deployment on Ascend NPUs. The 7B variant supports three distinct Chain-of-Thought (CoT) reasoning paradigms, namely slow_think, auto_think, and no_think, while the 1B variant operates exclusively in the no_think mode, which employs condensed reasoning for higher efficiency. Although CoT reasoning enhances capability, the generation of extended reasoning traces introduces substantial memory and latency overheads, posing challenges for practical deployment on Ascend NPUs. This paper addresses these computational constraints by leveraging low-bit quantization, which transforms FP16 computations into more efficient integer arithmetic. We introduce a unified low-bit inference framework, supporting INT8 (W8A8) and W4A8 quantization, specifically optimized for openPangu-Embedded models on the Atlas A2. Our comprehensive evaluation on code generation benchmarks (HumanEval and MBPP) demonstrates the efficacy of this approach. INT8 quantization consistently preserves over 90% of the FP16 baseline accuracy and achieves a 1.5x prefill speedup on the Atlas A2. Furthermore, W4A8 quantization significantly reduces memory consumption, albeit with a moderate trade-off in accuracy. These findings collectively indicate that low-bit quantization effectively facilitates efficient CoT reasoning on Ascend NPUs, maintaining high model fidelity.
@@ 1 Introduction
Large language models are increasingly being deployed on edge devices to facilitate on-device reasoning and code generation (Zheng et al., 2025; Girija et al., 2025). Recently, Huawei introduced the openPangu-Embedded series, which includes the 7B and 1B variants (Chen et al., 2025). The 7B variant supports three Chain-of-Thought (CoT) reasoning paradigms including slow_think, auto_think, and no_think (Wei et al., 2023a; Zhang et al., 2022), which differ in the verbosity and structure of reasoning steps, thereby influencing latency and output quality. The 1B variant implements only the no_think paradigm, producing condensed reasoning by default. Despite the strong performance of these models, their original FP16 versions require substantial memory and computational resources, which limit their practical deployment on NPU accelerators (Xiao et al., 2024; Ashkboos et al., 2024; Zhao et al., 2024), such as the Huawei Atlas A2, where memory bandwidth and storage capacity present significant constraints.
@@
To address the high memory and computational demands of FP16 inference, we apply quantization to the openPangu-Embedded models. This technique replaces floating-point parameters and activations with low-bit integer representations, significantly reducing both model size and computational complexity while aiming to maintain task performance (Wei et al., 2023b; Yao et al., 2023; Liu et al., 2024a). Specifically, we focus on two quantization configurations commonly used for efficient large language model deployment, namely INT8 (W8A8) and W4A8 (Dettmers et al., 2022; Lee et al., 2024b). Both configurations leverage the integer arithmetic capabilities of Ascend NPUs, offering a practical trade-off between efficiency and accuracy for on-device applications.
@@
However, most existing quantization techniques are designed and optimized for GPU architectures and are not easily transferable to Ascend NPUs(Wang et al., 2019). Adapting quantized models to the Atlas A2 presents significant challenges due to its unique compute primitives, memory hierarchy, and integer execution model. Efficient low-bit inference on this architecture requires tight integration between the quantization format, memory layout, and compute primitives to minimize redundant data movement and avoid underutilizing hardware resources (Feng et al., 2022; Frantar et al., 2023). To facilitate practical evaluation, we develop hardware-aware quantized inference paths that are fully optimized for the Atlas A2 execution environment.
@@
Beyond system-level deployment, we observe that quantization impacts the reasoning behavior of the openPangu-Embedded models. The three CoT paradigms slow_think, auto_think, and no_think exhibit distinct characteristics and respond differently to low-bit conversion(Chen et al., 2025). In slow_think, the model generates detailed intermediate steps to support complex reasoning. In no_think, the reasoning process is more condensed, retaining the stepwise structure but operating with higher efficiency. The auto_think mode adaptively switches between these two strategies based on input characteristics. Our evaluation reveals that quantization affects the behavior of all three modes, prompting a systematic investigation into the impact of quantization on both model execution and reasoning quality.
@@
(3) Analysis of how quantization affects CoT reasoning. We analyze the impact of INT8 quantization on CoT reasoning behavior in the openPangu-Embedded-7B across the three reasoning modes and in the 1B model under its native no_think mode, using HumanEval and MBPP as evaluation benchmarks. Our findings show that the transition from FP16 to INT8 introduces minimal disruption to the reasoning process in most scenarios, indicating that quantized models can still effectively perform structured multi-step inference.
@@ 2 Preliminary
To enable efficient inference of large language models on resource-constrained hardware, quantization techniques are used to approximate high-precision floating-point values with low-bit integer representations. This approach reduces both memory footprint and computational cost, while aiming to preserve model accuracy (Dettmers et al., 2022; Yao et al., 2023; Jin et al., 2024).
@@
The fidelity of quantization is highly dependent on the granularity of the scaling factor(Nagel et al., 2021). Per-tensor quantization uses a single scale for an entire tensor, offering memory efficiency but often resulting in lower accuracy(Bondarenko et al., 2021). Per-channel quantization assigns a separate scale to each output channel of weight matrices and is widely used due to its effectiveness (Lin et al., 2024b). For activations, per-token quantization computes one scale per token across its entire feature dimension (Shao et al., 2024). To balance flexibility and overhead, group-wise quantization divides each tensor into fixed-size groups, each with its own scale (Dettmers et al., 2023; Zhao et al., 2024). These strategies are essential for maintaining accuracy under low-bit constraints.
@@ 3.2 Quantization Configurations and Evaluation Protocol
The first variant adopts SmoothQuant (Xiao et al., 2024), which balances the quantization difficulty between activations and weights using a diagonal scaling ${S}=\text{diag}({s})$. The linear transformation is rewritten as:
@@
The second variant employs Hadamard rotation (Ashkboos et al., 2024; Liu et al., 2024b), which applies an orthogonal transform to the weight matrix before quantization. The output is computed as:
@@
where H{H} is a normalized Hadamard matrix. This rotation preserves the mathematical equivalence in full precision while promoting a more uniform weight distribution, which is better suited for low-bit quantization. This effect is visually demonstrated in Figure 1.
@@ 4.1 Experimental Setup
Quantization Configuration. We apply quantization without retraining, using calibrated scales derived from downstream task data. All quantized models employ symmetric quantization. Our evaluation covers INT8 and three W4A8 configurations, including a baseline W4A8 setting as well as two enhanced variants that incorporate SmoothQuant(Xiao et al., 2024) and Hadamard rotation(Ashkboos et al., 2024) respectively. For performance measurement on the Atlas A2, all configurations are deployed using hardware-optimized GEMM interfaces tailored to their corresponding bitwidths. Accuracy evaluation is conducted under a unified quantization pipeline to ensure consistent calibration and quantization procedures across all settings.
Benchmarks. We evaluate model performance on two standard code generation benchmarks, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). For each benchmark, we test all three CoT reasoning paradigms supported by openPangu-Embedded, including slow_think, auto_think, and no_think. We compare INT8 quantization against the FP16 baseline to assess accuracy retention and inference efficiency, and further analyze how the evaluated W4A8 configurations influence model fidelity under a fixed 4-bit weight setting on Ascend NPUs.
CoT Mode Activation. For the 7B model, slow_think is the default reasoning behavior and requires no modification to the input prompt. The auto_think and no_think modes are activated by appending corresponding directives (e.g., /auto_think; see Figure 2) to the end of the input prompt. In contrast, the 1B model operates exclusively in no_think mode by design. Its default reasoning behavior is inherently condensed and efficient, requiring no directive in the prompt.
Figure 2: Example of a user prompt under the auto_think mode. The reasoning directive no_think (highlighted in red) is appended to the prompt to suppress explicit CoT generation.
@@ 4.2 Main Results
Table 1 reports the accuracy of openPangu-Embedded models on HumanEval and MBPP under FP16 and INT8 quantization. For the 7B model, INT8 matches or exceeds the performance of FP16 across all settings. In no_think, the scores are similar (HumanEval: 85.37 vs 85.37; MBPP: 77.04 vs 78.21). In auto_think, INT8 improves by 2.44 on HumanEval and 3.11 on MBPP. Under slow_think, INT8 achieves 95.73 on HumanEval and 79.77 on MBPP, compared to FP16’s 95.12 and 77.43, respectively. For the 1B model, which only supports slow_think, INT8 slightly outperforms FP16 (HumanEval: 66.46 vs. 65.24; MBPP: 62.26 vs. 61.87). These results confirm that INT8 quantization preserves high code generation accuracy across model scales and reasoning strategies, supporting efficient deployment without significant loss in fidelity.
@@
1B no_think FP16 65.24 61.87
@@ 4.4 CoT Analysis
We analyze the impact of quantization on CoT output length across the openPangu-Embedded-7B model across its three supported reasoning modes, and compare with the 1B variant operating exclusively in no_think mode. As shown in Figure 3, quantization from FP16 to INT8 has little effect on output length in most configurations. Notably, the 7B model consistently produces shorter CoT traces than the 1B model in no_think mode, suggesting that larger models may adopt more concise reasoning strategies.
Figure 3: Average word count on HumanEval and MBPP for FP16 and INT8 openPangu-Embedded models.
@@
Beyond surface-level variations in output length and phrasing, we further examine the structural integrity of CoT reasoning under quantization. Specifically, we identify and analyze a critical degradation pattern termed repetitive generation, where model outputs terminate with identical phrases repeated multiple times. Figure 5 quantifies this phenomenon across configurations on HumanEval. The 1B model exhibits high susceptibility to repetitive outputs under FP16 (34.15%), which is substantially reduced to 21.95% under INT8 quantization. In contrast, the 7B model demonstrates inherent robustness across all three supported modes, with repetitive ratios consistently below 2.5% regardless of precision or mode. Crucially, repetitive generation strongly correlates with functional failure: non-repetitive samples achieve an average accuracy of 87.39%, while repetitive samples achieve only 18.24%. This indicates that repetitive generation fundamentally disrupts reasoning integrity.
Figure 5: Frequency of repetitive generation patterns across reasoning modes and model configurations on HumanEval. Repetitive generation is defined as terminal output segments containing identical phrases repeated until sequence termination. Bars represent the percentage of test samples exhibiting this behavior under FP16 and INT8 precision for 1B (no_think) and 7B (all three modes) model variants.
@@
Post-training quantization has become a prominent approach for compressing large language models without requiring access to original training data. It is broadly classified into weight-only quantization and weight–activation quantization. Weight-only approaches (Lin et al., 2024b; Frantar et al., 2023; Kim et al., 2024) compress model weights into low-bit representations while restoring them to high precision (e.g., FP16) during GEMM computations. Although this strategy alleviates memory bandwidth demands, the underlying arithmetic operations remain in high precision, which continues to impose a substantial bottleneck on inference efficiency (Lin et al., 2025). Consequently, significant opportunities remain for accelerating large language model inference. In contrast, weight–activation methods (Yao et al., 2022; Lee et al., 2024a) quantize both weights and activations into low-bit formats, enabling GEMM to be performed entirely in low precision. This paradigm mitigates both memory bandwidth and computational constraints but frequently incurs pronounced accuracy degradation due to outlier activations. To address this issue, mathematically equivalent transformation techniques (Xiao et al., 2024; Shao et al., 2024) employ channel-wise smoothing strategies. By transferring activation outliers into the weight domain, these methods effectively suppress quantization errors. More recently, rotation-based weight–activation methods (Lin et al., 2024a; Liu et al., 2024b; Ashkboos et al., 2024) have been proposed and demonstrated remarkable success in maintaining model accuracy even at 4-bit precision. Complementing uniform low-bit schemes, mixed-precision quantization achieves a better accuracy-efficiency trade-off by assigning higher bitwidths to outlier-prone components and lower bitwidths to stable ones (Zhao et al., 2024; Liu et al., 2025b).
Impact of quantization on reasoning behavior. Recent studies have shown that quantization can affect the reasoning behavior of large language models beyond mere task accuracy. Empirical evidence suggests that quantization error may accumulate over CoT steps, disproportionately degrading performance on complex reasoning tasks (Liu et al., 2025a). Furthermore, while moderate quantization typically preserves output length, which serves as a proxy for reasoning depth, aggressive low-bit schemes can lead to unstable generation and disrupted logical flow (Li et al., 2025). However, these analyses are primarily focused on general benchmarks and NVIDIA platforms, with limited attention given to structured reasoning modes like slow_think or auto_think on specialized hardware, such as Ascend NPUs.